# Parametric Scaling Law of Tuning Bias in Conformal Prediction

### v1

@@ 6 Related Works
Conformal prediction is a method for uncertainty quantification that ensures that the prediction intervals or sets cover the true label with a user-defined error rate (Papadopoulos, 2008; Vovk et al., 2005; Vovk, 2012). On the one hand, conditional coverage validity is a key property of conformal prediction, ensuring the algorithm fairness (Angelopoulos et al., 2021; Gibbs et al., 2024; Romano et al., 2020; Huang et al., 2024). On the other hand, to enhance the efficiency of conformal prediction, recent studies (Liu et al., 2025; Stutz et al., 2022; Kiyani et al., 2024b, a) have proposed training-based methods, which could be regarded as adapters, performing tuning prior to conformal prediction. Beyond these, with many non-conformity scores proposed, the score selection or aggregation is another challenge in conformal prediction (Yang & Kuchibhotla, 2024; Luo & Zhou, 2024; Fan et al., 2024; Gasparin & Ramdas, 2024; Ge et al., 2024; Qin et al., 2024). These recent paradigms of conformal prediction typically require tuning some parameters on a hold-out dataset, and our work provides the first study to quantify the negative effect induced by using the same dataset for tuning and calibration.
@@
As classical learnability theory, a learnable model could be regarded as a risk minimization model with a specific hypothesis class (Vapnik & Chervonenkis, 1971; Vapnik, 1991; van der Vaart & Wellner, 1996; Vapnik, 1999; Vershynin, 2018). And further, the constrained risk minimization model is developed as a special case of learnability theory, where the hypothesis class is the set with a specific constraint or structure, such as structural risk minimization of character recognition (Guyon et al., 1991), of data-dependent hypothesis class (Shawe-Taylor et al., 1998), of Rademacher penalty method (Koltchinskii, 2001) and rough set-based classifier (Liu et al., 2020). As for conformal prediction, several works study learn the conformal prediction directly by designing a specific objective function (Stutz et al., 2022; Noorani et al., 2024). Further, the conformal prediction is a special case of risk minimization problem with the constraint of coverage guarantee, and the generalization loss for the size of prediction set could be quantified (Gupta et al., 2022; Yang & Kuchibhotla, 2024), including classification (Bai et al., 2022) and regression (Gupta et al., 2022; Fan et al., 2024). As we proposed in Section 4, we interpret the tuning bias as a further constrained risk minimization problem of conformal prediction and provide the first theoretical analysis of the tuning bias.
@@ A.2 Setup for Parameter Tuning
given probability output $p:=p(y|\boldsymbol{x})$ for a classification task, and $u$ is a random variable drawn from the uniform distribution $U[0,1]$. and THR (Sadinle et al., 2019) defined as:
@@ B.1 Additional Results for Confidence Calibration
Figure 4: Additional results for confidence calibration across various datasets and score functions. The coverage gaps are obtained on the conformal prediction with a hold-out set or the same set as calibration set to parameter tuning. Figures 4(a) and 4(b) show the tuning bias with temperature scaling, using APS and THR as the score functions, respectively. The coverage gaps are obtained on the conformal prediction with a hold-out set or the same set as calibration set to parameter tuning. Figures 4(c) and 4(d) show the tuning bias with vector scaling, using THR as the score function. The experiments vary the size of the calibration set and the number of parameters. The experiments are conducted on pre-trained models in Section A.1. The coverage rate is set to 0.9. The figure shows that the tuning bias of temperature scaling is small for all datasets and score functions under various settings. The tuning bias of vector scaling is generally large except for CIFAR-10. For comparison, the tuning bias of temperature scaling is generally smaller than that of vector scaling.
@@ B.2 Additional Results for Score Function Tuning
We focus on two widely used score functions: RAPS and SAPS. The results for them are provided in Figures 6(a) and 6(b).
@@ B.3 Additional Results for Score Aggregation
In this study, we investigate the aggregation of three score functions—APS, RAPS, and SAPS—all derived from the same pre-trained classifier as a case example. The results are provided in Figure 6(c).
@@ B.4 Additional Results for C-Adapter
Additional experimental results on CovGap with C-Adapter are shown in Figures 6(d) and 6(e), using APS and THR as the score functions, respectively.
@@
Additional experimental results on CovGap with ConfTr (ft.) are shown in Figures 8(a) and 8(b), using APS and THR as the score functions, respectively.
@@
Appendix C Some Useful Lemmas and Corollaries
@@
Lemma C.2 (Gaussian Tail Inequality).
@@
=$\mathcal{H}=\{\mathds{1}\{S^{\lambda}(\boldsymbol{x},y)\leq t\}\mid\lambda\in\mathbb{R},t\in\mathbb{R}\}.$
If $S^{\lambda}(\boldsymbol{x},y)$ is continuous, bounded over $\lambda$ for any fixed $(\boldsymbol{x},y)$, then $\text{VC}(\mathcal{H})\leq 2$. Specifically, if $S^{\lambda}(\boldsymbol{x},y)$ is distinct over $\mathcal{X}\times\mathcal{Y}$ for any fixed $\lambda$, then $\text{VC}(\mathcal{H})=2$.
@@
We prove that $\text{VC}(\mathcal{H})\leq 2$ in two steps:
@@
    We show that $\text{VC}(\mathcal{H})\geq 2$. Given any two distinct points $(\boldsymbol{x}_{1},y_{1})$ and $(\boldsymbol{x}_{2},y_{2})$, we need to verify that all four possible binary labellings (0,0), (0,1), (1,0), (1,1) can be realized by some choice of $\lambda$ and $t$. Consider $S^{\lambda}(\boldsymbol{x},y)$:
@@
    Therefore, $\text{VC}(\mathcal{H})\geq 2$.
@@
    We prove that $\text{VC}(\mathcal{H})\leq 2$. Suppose for contradiction that three distinct points can be shattered. Consider the labelling pattern (1,0,1). This would require some $\lambda$ and $t$ such that:
@@
    implying $S^{\lambda}(\boldsymbol{x}_{2},y_{2})>\max\{S^{\lambda}(\boldsymbol{x}_{1},y_{1}),S^{\lambda}(\boldsymbol{x}_{3},y_{3})\}$. However, since $\lambda$ is one-dimensional and $S^{\lambda}(\boldsymbol{x},y)$ is continuous in $\lambda$, such a complex ordering cannot be achieved at a single $\lambda$ value for all possible three-point configurations. This phenomenon is because the continuity of $S^{\lambda}(\boldsymbol{x},y)$ over $\mathbb{R}$ ensures that any crossing of function values must occur in pairs, making it impossible to maintain the required ordering consistently. Therefore, three points cannot be shattered, and $\text{VC}(\mathcal{H})\leq 2$.
Combining both results, we conclude that $\text{VC}(\mathcal{H})=2$. ∎
Corollary C.4.
@@
=$\mathcal{H}=\{\mathds{1}\{S^{\lambda}(\boldsymbol{x},y)\leq t\}\mid\lambda\in\mathbb{R}^{+},t\in\mathbb{R}\}.$
If $S^{\lambda}(\boldsymbol{x},y)$ is continuous and bounded over $\lambda$ for any fixed $(\boldsymbol{x},y)$ and $S^{\lambda}(\boldsymbol{x},y)$ is distinct over $\mathcal{X}\times\mathcal{Y}$ for any fixed $\lambda$, then $\text{VC}(\mathcal{H})=2$.
@@
The proof mirrors Lemma C.3, with $\lambda>0$. This restriction does not alter the VC dimension: First, we show that $\text{VC}(\mathcal{H})\geq 2$. Two points can still achieve all labellings with positive $\lambda$ and $t$. Next, we show that $\text{VC}(\mathcal{H})\leq 2$. The restriction to $\lambda>0$ does not allow shattering three points. Thus, the VC dimension is 2. ∎
Lemma C.5.
@@
=$\mathcal{H}=\{\mathds{1}\{S^{\lambda}(x,y)\leq t\}\mid\lambda\in\mathbb{R}^{d},t\in\mathbb{R}\}.$
If $S^{\lambda}(x,y)$ is continuous and bounded over $\lambda$ for any fixed $(x,y)$, then $\text{VC}(\mathcal{H})\leq d+1$.
@@
We prove that $\text{VC}(\mathcal{H})\leq d+1$. Suppose for contradiction that $d+2$ distinct points can be shattered. Consider any labelling pattern that requires one point to have a value strictly larger than $d+1$ other points at the same $\lambda$. This would require some $\lambda$ and $t$ such that:
@@
where $|I|=d+1$. This implies $S^{\lambda}(\boldsymbol{x}_{j},y_{j})>\max_{i\in I}\{S^{\lambda}(\boldsymbol{x}_{i},y_{i})\}$. However, since $\lambda$ is $d$-dimensional and $S^{\lambda}(\boldsymbol{x},y)$ is continuous in $\lambda$, such a complex ordering requiring one point to dominate $d+1$ other points cannot be achieved at a single $\lambda$ value. This phenomenon is because the continuity of $S^{\lambda}(\boldsymbol{x},y)$ over $\mathbb{R}^{d}$ ensures that the ordering relationships between $d+2$ points cannot be arbitrary with only $d$ degrees of freedom. Therefore, $d+2$ points cannot be shattered, and $\text{VC}(\mathcal{H})\leq d+1$. ∎
Lemma C.6 (Empirical Process Bound via VC Dimension).
For the class $\mathcal{H}$ with finite VC dimension $\text{VC}(\mathcal{H})\geq 1$, we have
$\mathbb{E}\mathfrak{R}_{\Lambda,\mathcal{D}_{\text{cal}}}\leq C\sqrt{\frac{\text{VC}(\mathcal{H})}{n}},$
@@
Appendix D Proof of Theorem 4.1
@@
$\mathop{CovGap}(\widehat{\mathcal{C}},\alpha)\leq\mathbb{E}\mathfrak{R}_{\Lambda,\mathcal{D}}+\varepsilon_{\alpha,n}.$
Appendix E Proof of Proposition 4.2
@@
where the second inequality is due to Lemma C.2. ∎
Appendix F Proof of Corollary 4.3 and 4.4
@@
Appendix G Proof of Proposition 4.6
@@
It is a direct application of Lemma C.5 and C.6. ∎
Appendix H Proof of Corollary 4.7
@@
Appendix I Proof of Proposition 5.1
@@
Then, for any two samples $(\boldsymbol{x},y)$ and $(\boldsymbol{x}^{\prime},y^{\prime})$, with the property of softmax function $\psi_{y}$ on binary classification (see Lemma C.8 in Appendix C), we have
@@
Therefore, the order of the scores on the dataset $\mathcal{D}_{\text{cal}}$ is preserved for any transformation $\lambda\in\mathbb{R}^{+}$, i.e, the order of $\mathcal{S}^{\lambda}_{\mathcal{D}_{\text{cal}}}$ is the same as $\mathcal{S}_{\mathcal{D}_{\text{cal}}}$ for any $\lambda\in\mathbb{R}^{+}$. Then by Lemma C.7 in Appendix C, we have
@@
for any $\lambda\in\mathbb{R}^{+}$, which means $\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{TS}})=0\leq\mathop{TuningBias}(\widehat{\mathcal{C}}_{\text{VS}})$.
Appendix J Order-preserving Regularization and Proof of Proposition 5.2
@@
$\mathcal{H}_{\text{op}}=\{f_{\lambda,t}\in\mathcal{H}\text{ s.t. }o(y,\boldsymbol{x},p)=o(y,\boldsymbol{x},p_{\lambda})\},$ (14)
where $o(y,\boldsymbol{x},p)$ is the order of the true label $y$ in the sorted predicted probabilities $p(\cdot|\boldsymbol{x})$, and $p_{\lambda}$ is the probability after confidence calibration like temperature scaling or vector scaling. The class $\mathcal{H}_{\text{op}}$ is a constraint version of $\mathcal{H}$ with a special regularization. With the constrained ERM with artificial order-preserving regularization, we could obtain a less tuning bias.

### v2

@@ 6 Related Works
Conformal prediction is a method for uncertainty quantification that ensures that the prediction intervals or sets cover the true label with a user-defined error rate (Papadopoulos, 2008; Vovk et al., 2005; Vovk, 2012). On the one hand, conditional coverage validity is a key property of conformal prediction, ensuring the algorithm fairness (Angelopoulos et al., 2021; Gibbs et al., 2024; Romano et al., 2020; Huang et al., 2024). On the other hand, to enhance the efficiency of conformal prediction, recent studies (Liu et al., 2025; Stutz et al., 2022; Kiyani et al., 2024b, a) have proposed training-based methods, which could be regarded as adapters, performing tuning prior to conformal prediction. Beyond these, with many non-conformity scores proposed, the score selection or aggregation is another challenge in conformal prediction (Yang & Kuchibhotla, 2024; Luo & Zhou, 2024; Fan et al., 2024; Gasparin & Ramdas, 2024; Ge et al., 2024; Qin et al., 2024). These recent paradigms of conformal prediction typically require tuning parameters on a hold-out dataset, and our work provides the first study to quantify the negative effect induced by using the same dataset for tuning and calibration.
@@
As classical learnability theory, a learnable model could be regarded as a risk minimization model with a specific hypothesis class (Vapnik & Chervonenkis, 1971; Vapnik, 1991; van der Vaart & Wellner, 1996; Vapnik, 1999; Vershynin, 2018). And further, the constrained ERM is developed as a special case of learnability theory, where the hypothesis class is the set with a specific constraint or structure, such as structural risk minimization of character recognition (Guyon et al., 1991), of data-dependent hypothesis class (Shawe-Taylor et al., 1998), of Rademacher penalty method (Koltchinskii, 2001) and rough set-based classifier (Liu et al., 2020). As for conformal prediction, several works study learn the conformal prediction directly by designing a specific objective function (Stutz et al., 2022; Noorani et al., 2024). Further, the conformal prediction is a special case of risk minimization problem with the constraint of coverage guarantee, and the generalization loss for the size of prediction set could be quantified (Gupta et al., 2022; Yang & Kuchibhotla, 2024), including classification (Bai et al., 2022) and regression (Gupta et al., 2022; Fan et al., 2024). As we proposed in Section 4, we interpret the tuning bias as a further constrained risk minimization problem of conformal prediction and provide the first systematic theoretical analysis of the tuning bias, offering a novel tool for analyzing the learnability of general conformal prediction problems.
@@ 7 Conclusion
Acknowledgement
We thank the anonymous reviewers for their valuable comments and suggestions. This research is supported by the Shenzhen Fundamental Research Program (Grant No.JCYJ20230807091809020), the SUSTech-NUS Joint Research Program and in part by the National Natural Science Foundation of China (NSFC) under grant 12371290. We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology.
@@ A.2 Setup for Parameter Tuning
given probability output $p:=p(y|\boldsymbol{x})$ for a classification task, and $u$ is a random variable drawn from the uniform distribution $U[0,1]$, and THR (Sadinle et al., 2019) defined as:
@@ Appendix B Additional Experimental Results
By the results in this section, it is worthy noting that the parameter numbers of those tuning methods (e.g., VS and C-Adapter) are positively related to the class numbers of the dataset. Thus, datasets with more classes require more parameter numbers in the tuning, leading to a larger tuning bias. This explains why those methods perform differently in various datasets. The details are provided as following.
@@ B.1 Additional Results for Confidence Calibration
The results in Figure 4 show that the tuning bias of temperature scaling is small for all datasets and score functions under various settings. The tuning bias of vector scaling is generally large except for CIFAR-10. For comparison, the tuning bias of temperature scaling is generally smaller than that of vector scaling.
@@
Figure 4: Additional results for confidence calibration across various datasets and score functions. The coverage gaps are obtained on the conformal prediction with a hold-out set or the same set as calibration set to parameter tuning. Figures 4(a) and 4(b) show the tuning bias with temperature scaling, using APS and THR as the score functions, respectively. The coverage gaps are obtained on the conformal prediction with a hold-out set or the same set as calibration set to parameter tuning. Figures 4(c) and 4(d) show the tuning bias with vector scaling, using THR as the score function. The experiments vary the size of the calibration set and the number of parameters. The experiments are conducted on pre-trained models in Section A.1. The coverage rate is set to 0.9. The figure shows that the tuning bias of temperature scaling is small for all datasets and score functions under various settings. The tuning bias of vector scaling is generally large except CIFAR-10. For comparison, the tuning bias of temperature scaling is generally smaller than that of vector scaling.
@@ B.2 Additional Results for Score Function Tuning
We focus on two widely used score functions: RAPS and SAPS. The results for them are provided in Figures 6(a) and 6(b). The results in these figures shows the tunig bias of RAPS and SAPS is not significantly different from the one with a hold-out set under various settings.
@@ B.3 Additional Results for Score Aggregation
In this study, we investigate the aggregation of three score functions—APS, RAPS, and SAPS—all derived from the same pre-trained classifier as a case example. The results are provided in Figure 6(c). The results in this figure shows the tuning bias of score aggregation is not significantly different from the one with a hold-out set under various settings.
@@ B.4 Additional Results for C-Adapter
Additional experimental results on CovGap with C-Adapter are shown in Figures 6(d) and 6(e), using APS and THR as the score functions, respectively. The results in these figures shows the tuning bias of C-Adapter is not significantly different from the one with a hold-out set under various settings.
@@
Additional experimental results on CovGap with ConfTr (ft.) are shown in Figures 8(a) and 8(b), using APS and THR as the score functions, respectively. The results in these figures shows the tuning bias of ConfTr (ft.) is significantly different from the one with a hold-out set under various settings except CIFAR-10.
@@
Appendix C Additional Results for Regression
We clarify that our analysis reveals a general phenomenon of conformal prediction in both classification and regression tasks. Section 3 focuses on classification, as many parameter tuning methods (like TS/VS and ConfTr) are designed for classification tasks. Here, we provide the results for regressions to show the case of regression tasks following previous work (Liang et al., 2024) with a target coverage of 90% and 30 repetitions. The details are presented as following:
We conduct an empirical study to evaluate methods for constructing prediction intervals using the Conformalized Quantile Regression (CQR) framework. The experiments are performed on the Protein Structure dataset obtained from the UCI repository, which comprises $N=45730$ data points with feature dimension $d=9$. The base learners employed are RandomForestQuantileRegressor instances from Python’s quantile_forest package. The target nominal coverage level for all experiments is set to $0.91-\alpha=$ 0.9.
We investigate two distinct strategies for utilizing the data reserved after model training (referred to as calibration set and hold-out set): Same: A single dataset is allocated for both model selection (i.e., choosing the optimal hyperparameter configuration from a set of candidates) and the subsequent calibration step (i.e., determining the empirical quantile required to adjust the prediction interval width); Hold-out: The reserved data is partitioned into two disjoint sets of equal size, where the first set is used for model selection, while the second set is used for calibration, consequently requiring twice the amount of reserved data compared to the Shared strategy to achieve an equivalently sized dataset for the final calibration phase. Here we consider the following two experiments:
Experiment 1: Varying the number of candidate models.
For the first experiment, the size of calibration set in the Shared strategy is fixed at $n=50$. For the Shared strategy, these $50$ data points are employed for both model selection and calibration, while for the Split strategy, $50$ data points are used for model selection, and a separate set of $50$ data points is used for calibration, amounting to $100$ reserved data points in total. The complexity of the model selection task is modeled by varying the number of candidate models. All candidate models are RandomForestQuantileRegressor instances. The max_features parameter is varied across 10 distinct values, specifically np.linspace(d/10, d, 10). The n_estimator parameter is adjusted to alter the total pool of models: We vary the number of candidate models by adjusting the n_estimator parameter to 4, 8, 12, 16, and 20 values, which combined with the 10 distinct max_features values results in 40, 80, 120, 160, and 200 candidate models respectively.
Experiment 2: Varying the Calibration Set Size
In this experiment, the number of candidate models is held constant at 200 (derived from 20 n_estimator values and 10 max_features values). We vary the size of the calibration dataset, denoted as $n$. For the Same strategy, $n$ represents the size of the single dataset used for both model selection and calibration, while for the Hold-out strategy, $n$ denotes the size of each of the two datasets (one for selection and one for calibration), resulting in a total reserved data of $2\times n$ for the Hold-out strategy. The values investigated for both strategies are $n\in\{100,200,300,400,500\}$.
The results of the two experiments are summarized in Table 2 and Table 3, respectively. Mean Coverage and Std Coverage are calculated from the 30 independent runs. The results of the two experiments are similar to the ones in the main text.
Table 2: Results for varying the number of candidate models. For the Same strategy, calibration size $n=50$. For the Hold-out strategy, $n=50$ for model selection and $n=50$ for calibration. Target coverage $0.91-\alpha=$ 0.9.
# Models Method Coverage Length Coverage Gap Tuning Bias
Mean (Std) Mean (Std) Mean (Std) Mean
40 Same 0.8437 (0.0383) 12.89 (1.22) 0.0563 (0.0348) 0.0246
Hold-out 0.9251 (0.0268) 17.28 (2.96) 0.0317 (0.0212) -
80 Same 0.8229 (0.0354) 12.41 (1.34) 0.0771 (0.0353) 0.0412
Hold-out 0.9049 (0.0404) 16.01 (2.51) 0.0359 (0.0252) -
120 Same 0.8258 (0.0564) 12.27 (1.63) 0.0742 (0.0520) 0.0324
Hold-out 0.8953 (0.0480) 15.94 (3.17) 0.0418 (0.0293) -
160 Same 0.8373 (0.0467) 12.40 (1.33) 0.0627 (0.0413) 0.0371
Hold-out 0.9163 (0.0286) 16.24 (2.06) 0.0256 (0.0217) -
200 Same 0.8447 (0.0494) 12.44 (1.41) 0.0553 (0.0429) 0.0187
Hold-out 0.9037 (0.0406) 15.09 (2.25) 0.0366 (0.0246) -
Table 3: Results for varying the calibration set size ($n$). The number of candidate models is fixed at 200. For the Hold-out strategy, $n$ refers to the size of each of the two sets (selection and calibration), totaling $2n$ reserved points. Target coverage $0.91-\alpha=$ 0.9.
$n$ Method Coverage Length Coverage Gap Tuning Bias
Mean (Std) Mean (Std) Mean (Std) Mean
100 Same 0.8459 (0.0509) 12.77 (1.25) 0.0541 (0.0433) 0.0150
Hold-out 0.9079 (0.0456) 16.16 (3.11) 0.0391 (0.0278) -
200 Same 0.8553 (0.0378) 13.10 (0.96) 0.0448 (0.0337) 0.0118
Hold-out 0.9043 (0.0394) 15.12 (1.72) 0.0330 (0.0212) -
300 Same 0.8608 (0.0326) 13.04 (0.93) 0.0392 (0.0296) 0.0158
Hold-out 0.8967 (0.0323) 14.46 (1.04) 0.0234 (0.0192) -
400 Same 0.8595 (0.0292) 13.13 (0.56) 0.0405 (0.0266) 0.0157
Hold-out 0.8927 (0.0322) 14.37 (1.16) 0.0248 (0.0201) -
500 Same 0.8767 (0.0388) 13.25 (0.68) 0.0327 (0.0269) 0.0037
Hold-out 0.9007 (0.0323) 14.21 (0.65) 0.0290 (0.0189) -
Appendix D Some Useful Lemmas and Corollaries
@@
Lemma D.2 (Gaussian Tail Inequality).
@@
$\mathcal{H}_{\Lambda}=\{\mathds{1}\{S^{\lambda}(\boldsymbol{x},y)\leq t\}\mid\lambda\in\mathbb{R},t\in\mathbb{R}\}.$
If $S^{\lambda}(\boldsymbol{x},y)$ is continuous, bounded over $\lambda$ for any fixed $(\boldsymbol{x},y)$, then $\text{VC}(\mathcal{H}_{\Lambda})\leq 2$. Specifically, if $S^{\lambda}(\boldsymbol{x},y)$ is distinct over $\mathcal{X}\times\mathcal{Y}$ for any fixed $\lambda$, then $\text{VC}(\mathcal{H}_{\Lambda})=2$.
@@
We prove that $\text{VC}(\mathcal{H}_{\Lambda})\leq 2$ in two steps:
@@
    We show that $\text{VC}(\mathcal{H}_{\Lambda})\geq 2$. Given any two distinct points $(\boldsymbol{x}_{1},y_{1})$ and $(\boldsymbol{x}_{2},y_{2})$, we need to verify that all four possible binary labellings (0,0), (0,1), (1,0), (1,1) can be realized by some choice of $\lambda$ and $t$. Consider $S^{\lambda}(\boldsymbol{x},y)$:
@@
    Therefore, $\text{VC}(\mathcal{H}_{\Lambda})\geq 2$.
@@
    We prove that $\text{VC}(\mathcal{H}_{\Lambda})\leq 2$. Suppose for contradiction that three distinct points can be shattered. Consider the labelling pattern (1,0,1). This would require some $\lambda$ and $t$ such that:
@@
    implying $S^{\lambda}(\boldsymbol{x}_{2},y_{2})>\max\{S^{\lambda}(\boldsymbol{x}_{1},y_{1}),S^{\lambda}(\boldsymbol{x}_{3},y_{3})\}$. However, since $\lambda$ is one-dimensional and $S^{\lambda}(\boldsymbol{x},y)$ is continuous in $\lambda$, such a complex ordering cannot be achieved at a single $\lambda$ value for all possible three-point configurations. This phenomenon is because the continuity of $S^{\lambda}(\boldsymbol{x},y)$ over $\mathbb{R}$ ensures that any crossing of function values must occur in pairs, making it impossible to maintain the required ordering consistently. Therefore, three points cannot be shattered, and $\text{VC}(\mathcal{H}_{\Lambda})\leq 2$.
Combining both results, we conclude that $\text{VC}(\mathcal{H}_{\Lambda})=2$. ∎
Corollary D.4.
@@
$\mathcal{H}_{\Lambda}=\{\mathds{1}\{S^{\lambda}(\boldsymbol{x},y)\leq t\}\mid\lambda\in\mathbb{R}^{+},t\in\mathbb{R}\}.$
If $S^{\lambda}(\boldsymbol{x},y)$ is continuous and bounded over $\lambda$ for any fixed $(\boldsymbol{x},y)$ and $S^{\lambda}(\boldsymbol{x},y)$ is distinct over $\mathcal{X}\times\mathcal{Y}$ for any fixed $\lambda$, then $\text{VC}(\mathcal{H}_{\Lambda})=2$.
@@
The proof mirrors Lemma D.3, with $\lambda>0$. This restriction does not alter the VC dimension: First, we show that $\text{VC}(\mathcal{H}_{\Lambda})\geq 2$. Two points can still achieve all labellings with positive $\lambda$ and $t$. Next, we show that $\text{VC}(\mathcal{H}_{\Lambda})\leq 2$. The restriction to $\lambda>0$ does not allow shattering three points. Thus, the VC dimension is 2. ∎
Lemma D.5.
@@
$\mathcal{H}_{\Lambda}=\{\mathds{1}\{S^{\lambda}(x,y)\leq t\}\mid\lambda\in\mathbb{R}^{d},t\in\mathbb{R}\}.$
If $S^{\lambda}(x,y)$ is continuous and bounded over $\lambda$ for any fixed $(x,y)$, then $\text{VC}(\mathcal{H}_{\Lambda})\leq d+1$.
@@
We prove that $\text{VC}(\mathcal{H}_{\Lambda})\leq d+1$. Suppose for contradiction that $d+2$ distinct points can be shattered. Consider any labelling pattern that requires one point to have a value strictly larger than $d+1$ other points at the same $\lambda$. This would require some $\lambda$ and $t$ such that:
@@
where $|I|=d+1$. This implies $S^{\lambda}(\boldsymbol{x}_{j},y_{j})>\max_{i\in I}\{S^{\lambda}(\boldsymbol{x}_{i},y_{i})\}$. However, since $\lambda$ is $d$-dimensional and $S^{\lambda}(\boldsymbol{x},y)$ is continuous in $\lambda$, such a complex ordering requiring one point to dominate $d+1$ other points cannot be achieved at a single $\lambda$ value. This phenomenon is because the continuity of $S^{\lambda}(\boldsymbol{x},y)$ over $\mathbb{R}^{d}$ ensures that the ordering relationships between $d+2$ points cannot be arbitrary with only $d$ degrees of freedom. Therefore, $d+2$ points cannot be shattered, and $\text{VC}(\mathcal{H}_{\Lambda})\leq d+1$. ∎
Lemma D.6 (Empirical Process Bound via VC Dimension).
For the class $\mathcal{H}_{\Lambda}$ with finite VC dimension $\text{VC}(\mathcal{H}_{\Lambda})\geq 1$, we have
$\mathbb{E}\mathfrak{R}_{\Lambda,\mathcal{D}_{\text{cal}}}\leq C\sqrt{\frac{\text{VC}(\mathcal{H}_{\Lambda})}{n}},$
@@
Appendix E Proof of Theorem 4.1
@@
$\mathop{\text{CovGap}}({\mathcal{C}},\alpha)\leq\mathbb{E}\mathfrak{R}_{\Lambda,\mathcal{D}}+\varepsilon_{\alpha,n}.$
Appendix F Proof of Proposition 4.2
@@
where the second inequality is due to Lemma D.2. ∎
Appendix G Proof of Corollary 4.3 and 4.4
@@
Appendix H Proof of Proposition 4.6
@@
It is a direct application of Lemma D.5 and D.6. ∎
Appendix I Proof of Corollary 4.7
@@
Appendix J Proof of Proposition 5.1
@@
Then, for any two samples $(\boldsymbol{x},y)$ and $(\boldsymbol{x}^{\prime},y^{\prime})$, with the property of softmax function $\psi_{y}$ on binary classification (see Lemma D.8 in Appendix D), we have
@@
Therefore, the order of the scores on the dataset $\mathcal{D}_{\text{cal}}$ is preserved for any transformation $\lambda\in\mathbb{R}^{+}$, i.e, the order of $\mathcal{S}^{\lambda}_{\mathcal{D}_{\text{cal}}}$ is the same as $\mathcal{S}_{\mathcal{D}_{\text{cal}}}$ for any $\lambda\in\mathbb{R}^{+}$. Then by Lemma D.7 in Appendix D, we have
@@
for any $\lambda\in\mathbb{R}^{+}$, which means $\mathop{\text{TuningBias}}(\widehat{\mathcal{C}}_{\text{TS}})=0\leq\mathop{\text{TuningBias}}(\widehat{\mathcal{C}}_{\text{VS}})$.
Appendix K Order-preserving Regularization and Proof of Proposition 5.2
@@
$\mathcal{H}_{\lambda,\text{op}}=\{f_{\lambda,t}\in\mathcal{H}_{\Lambda}\text{ s.t. }o(y,\boldsymbol{x},p)=o(y,\boldsymbol{x},p_{\lambda})\},$ (14)
where $o(y,\boldsymbol{x},p)$ is the order of the true label $y$ in the sorted predicted probabilities $p(\cdot|\boldsymbol{x})$, and $p_{\lambda}$ is the probability after confidence calibration like temperature scaling or vector scaling. The class $\mathcal{H}_{\Lambda,\text{op}}$ is a constraint version of $\mathcal{H}_{\Lambda}$ with a special regularization. With the constrained ERM with artificial order-preserving regularization, we could obtain a less tuning bias. The tuning bias of the conformal prediction with order-preserving regularization is bounded by the tuning bias of the conformal prediction without order-preserving regularization as Proposition 5.2.